OpenStudy (anonymous):

Which expression is equivalent to 3x sqrt 3

OpenStudy (jadeishere):

Equivalent to what?

OpenStudy (anonymous):

3x sqrt 3

OpenStudy (jadeishere):

so \[3x \sqrt{3}\]

OpenStudy (anonymous):

ok now wat

OpenStudy (jadeishere):

Okay you're slightly confusing me.. What does the expression look like in your lesson, or whatever you're doing?

OpenStudy (anonymous):

just like youv typed it

OpenStudy (jadeishere):

Okay then, sorry I was confused at first. I believe x multiplies by sqrt3 works.. \[x \times \sqrt{3}\] @hartnn

hartnn (hartnn):

what are your options/choices ?

OpenStudy (the_fizicx99):

3x * (3)^1/2 ?